# Quark-versus-gluon tagging in CMS Open Data with CWoLa and TopicFlow

### v1

@@ IV.3 Classification
In addition to neural-network classifiers, we consider discriminating observables used in previous CMS studies [86, 87]. We consider the multiplicity of the jet constituents, the jet mass and the jet fragmentation distribution
@@
Finally, we also consider $\sigma_{1}$, the major axis of the jet determined by calculating the eigenvalues of the matrix
@@
In Fig. 8, we present both MC truth-derived (8a) and data-derived (8b, 8c, 8d) SIC curves for each classifier, shown as solid lines, and the aforementioned substructure observables, shown as dashed lines. Here, the label “TopicFlow” refers to a generative classifier of the form $p_{Q}(x)/p_{G}(x)$. The three data-derived panels assume different $f_{\mathrm{DoubleMu}},f_{\mathrm{Jet}}$ ($S$, $T$, $R$) as given in Tab. 4 . For the neural-network classifiers, curves are produced by averaging predictions of 10 independent models on the test set of the relevant dataset. Curves for the substructure observables are measured over the full dataset. The corresponding AUC scores are summarised in Tab. 5.
Figure 9: SIC curves constructed using samples from TopicFlow quark and gluon models. Each panel differs by the assumed quark fractions $f_{\mathrm{DoubleMu}},f_{\mathrm{Jet}}$ used to train the TopicFlow model. Error bands show the standard deviation across 100 (10$\times$10) TopicFlow/classifier pairs.
When evaluating classifiers on MC labels, as for Fig. 8a, the fully supervised classifier is best, closely followed by the MC-CWoLa classifier. Save for jet mass, all of the substructure observables outperform the classifiers trained on data (Data CWoLa and TopicFlow). Considering next the data SIC curves of Figs 8b, 8c and 8d, the rankings are different. In particular, the Data CWoLa classifier achieves the maximum significance improvement while the fully-supervised network is bested by the $\sigma_{1}$ observable. [^8] Note that the peaks at low $\varepsilon_{Q}$ are likely artefacts due to low statistics in the testing set. TopicFlow and MC CWoLa achieve similar results to one another and have the next best performance after Data CWoLa. Comparing these data-derived panels to one another, we see that the assumed quark fractions $f_{\mathrm{DoubleMu}}$ and $f_{\mathrm{Jet}}$ do not impact the rankings among the classifiers. What does change between these two constructions is the absolute significance improvement of the models as well as the signal efficiency at which it is achieved. Specifically, the maximum SIC is highest in Fig. 8c which uses fractions that assume quarks and gluons are mutually irreducible in Data CWoLa predictions ($T$). On the other hand, the SIC curves in Fig. 8c, which use fractions extracted by Data CWoLa assuming the same $\kappa_{QG}$ as MC, have the lowest maxima. In fact, both panels in the right column indicate that little-to-no significance improvement can be yielded by any cut on the fully-supervised networks. Notably, no classifier in any of the data-derived panels reaches the maximum SIC achieved by the fully supervised classifier in MC (Fig. 8a).
Finally, Fig. 9 shows SIC curves of each discriminative classifier as estimated using samples from TopicFlow quark and gluon distributions. [^9] TopicFlow generative classifiers are excluded since they represent the precise likelihood ratio for their samples and are therefore biased. We show three panels corresponding to choices $S$, $T$ or $R$ for the fractions $f_{\mathrm{DoubleMu}}$ and $f_{\mathrm{Jet}}$. The error bands show one standard deviation across the collection of curves produced by evaluating 10 classifiers on samples from 10 TopicFlow models. These bands therefore capture uncertainty in the TopicFlow model, which we find comprises most of the error. The general shapes of the curves agree with Fig. 8 except at low $\varepsilon_{Q}$ where the SIC for Data CWoLa previously exhibited large statistical variations. When amplifying the statistics with TopicFlow generated samples, the curves for each classifier are smooth and clearly exhibit a maximum. The TopicFlow SIC curves also maintain the same hierarchy among the three classifiers as Fig. 8. We note that the width of the bands on each curve are smaller than the differences between panels in Figs. 9b and 9c. Thus, the largest source of uncertainty in the expected discrimination power for a given quark gluon tagger in our datasets comes from the definition of the classes themselves. Again, as one expects, the greatest peak SIC is predicted by assuming mutually irreducible quarks and gluons since these are by definition maximally-separable categories.
V Conclusions
Machine-learning methods that operate directly on real data offer a way to avoid test-time performance loss caused by imprecision of Monte Carlo simulation. In this work we have studied the application of weakly-supervised neural networks to quark/gluon classification using jets in the CMS Open Data. In particular, we took $Z+$jet and dijet events as respective quark- and gluon- enriched samples to train CWoLa and TopicFlow models. The performance of these models was compared to fully-supervised networks trained on MC simulation.
In addition to the predictions based on parton labels in simulation, we made two estimates for the quark fractions of each real dataset. The first assumes that quarks and gluons are irreducible in the output of the classifiers, in line with the jet-topic framework. The second estimate assumes non-zero reducibility between quarks and gluons under the classifiers, as determined by MC. The results of both estimates differ from the MC prediction.
Using the extracted fractions, TopicFlow networks were trained to model the underlying quark and gluon distributions in real data. We employed a variant of the loss function introduced in Ref. [42] that allows for more stable training due to a convex optimisation objective. The distributions encoded by these TopicFlow models fit the corresponding topics in data better than predictions from simulation—both pure quark/gluon categories and topics constructed in MC.
An interesting direction for future work would be to relax the assumption of sample independence of quarks and gluons in the mixed datasets. While we observed that sample independence is approximately satisfied in the Monte Carlo, we did not attempt to quantify effects of its violation. As such, our results are subject to this extra degree of uncertainty. Some ideas to this end are given in Ref. [34]. A different approach may be to match the two mixed distributions via reweighting as in Ref. [9] (and implicitly in Ref. [31]). Alternatively, techniques that directly morph between two distributions could be used, such as in Refs. [83, 88, 89, 90]. Beyond cases where sample dependence is only approximately satisfied, a dedicated treatment of this problem could extend the reach of weakly-supervised approaches to domains where the effect is known to be severe [91].
Appendix A The CMS Open Data
A.1 Accessing the Open Data
The data from the 2010–2012 period can be accessed remotely or downloaded in the Analysis Object Data (AOD) format, a ROOT-compatible collection of data processed through the CMS software framework CMSSW [92]. The AODs are grouped into primary datasets on the basis of High-Level Trigger (HLT) paths that share common physics content. The typical size of an AOD file is a few gigabytes, and some thousands of AOD files constitute a primary dataset, which typically has a total size on the order of terabytes. Since we are interested in sourcing data samples with enriched quark- and gluon-jet content, our study makes use of the Jet [59] and DoubleMu [75] primary datasets, from which we extract gluon-rich dijet events and quark-rich $Z$ + jets events, respectively.
Our study of the 2011 Open Data takes place within version 5.3.32 of CMSSW, the recommended version compatible with the 2011 data. This software framework is used in data taking, processing, analysis and the generation of simulated samples. It defines the event model used to store the data and contains services and plugins we use in the initial stages of our data processing. The code is open source and available on GitHub [92].
The CMSSW framework is also available through the CernVM file system (CVMFS) [93]. The open data release is accompanied by a virtual machine (VM) image, to which CVMFS is mounted, and this provides a simple interface to access the data and conduct analysis. Rather than using the VM, we instead use a light-weight container image executed on a batch computing system. The image contains the base operating system needed to execute CMSSW. This, along with a CVMFS mount, allows us to run the CMS software framework with only the data and some additional resources we discuss below.
Our approach to accessing the data is to download each AOD file on demand through the XRootD protocol [94], pass the events through our selection and analysis code within CMSSW, and write out much smaller, custom data files in the HDF5 format for downstream processing. This workflow allows us to work with multiple primary datasets without keeping additional terabytes of information we do not require in our study.
Our experience using the CMS Open Data has greatly benefited from the detailed former studies using the CMS Open Data and the increasing number of example analyses provided by CMS through the cms-opendata-validation repositories [95]. Specifically, we have drawn on example code for the production of flat jet-tuples using the 2011 data [96] and the general procedures of Ref. [45] and the MODProducer [97], adapted to work with the prescaled dijet triggers of the 2011 Jet Primary Dataset. We point the interested reader to Refs. [45, 98, 99] for a detailed discussion of challenges and advice for using the CMS Open Data.
A.2 Analysing the Open Data
In addition to the AOD files, our analysis code requires access to the CMS condition database [100, 101]. This is a store of auxiliary, non-event-related data provided in the form of SQLite files within a local /cvmfs/cms-opendata-conddb.cern.ch directory inside the container image. In the context of our analysis these provide access to trigger information and to the jet energy corrections (JECs): MC-derived correction factors that mitigate detector noise and pile up. The condition database contains information relevant for different data-taking periods, each labelled by a global tag. Our analysis uses the global tags START53_LV6A1::All for simulation and FT_53_LV5_AN1::All for data. A connection is established to the conditions database within our configuration file.
The jet constituents are the objects obtained from the Particle Flow (PF) algorithm [102, 103]. In our analysis code, we pull jets from the ak5PFJets collection, corresponding to Fastjet anti-$k_{t}$ jets reconstructed from the PF particles with $R=0.5$. The constituent and jet $p_{T}$, along with the jet energy fractions (used later to implement our jet-quality requirements), are scaled by the JECs mentioned earlier. These are sourced using the ak5PFL1FastL2L3Residual correction label for data, and ak5PFL1FastL2L3 for MC.
Another layer of processing is required to deal with pileup: contamination from additional, non-event vertices. Here, we apply the procedure of Charged Hadron Subtraction (CHS) [104], where tracks not associated with the event vertex (defined below) are pruned from a jet. We draw vertex information from the standard vertex collection provided by CMSSW within our configuration file with the requirements that the minimum number of degrees of freedom in the vertex fit be larger than 4; $|z|\leq 24$ mm, where $z$ is the $z$-coordinate of the point of closest approach of the tracks to the $z$-axis; and $\rho<2$, where $\rho$ is an estimate of the average amount of transverse energy in the event per unit area coming from soft activity such as pileup. From among these vertices, the event vertex is defined as that primary vertex whose tracks have the largest sum of squared $p_{T}$, while all other interaction vertices are classed as pileup vertices. Within the EDAnalyzer we write out the vertex-fit information associated with the charged hadrons and later remove those assigned to pileup vertices. The distribution of the number of offline primary vertices (NPV) looks significantly different in simulation compared to data for the triggers of interest to our study. To correct this, we apply an additional weight to the MC that makes the NPV distributions match.
The remaining PF candidates are used to calculate the jet observables that form the basis of our machine-learning study. (These are introduced in more detail in Sec. III.) These observables are IRC-safe, and therefore an additional cut on the $p_{T}$ of the jet-constituent objects is unnecessary.
The muon objects that enter our analysis are required to pass the GlobalMuonPromptTight requirements. These include that the track should be identified as a global muon [77], the associated normalised $\chi^{2}$ in the global muon fit should be less than 10, and the number of valid muon-detector hits used in the fit should be non-zero.
In order to associate truth-level parton information with the PF candidates in MC, the generator objects must be matched to the PF objects from offline reconstruction (RECO) through MC matching. We follow the implementation of this matching from the CMSSW Physics Analysis Toolkit [105]. Specifically, we call the MCMatcher EDProducer within our configuration file. This matches jets from the ak5PFJets collection we use in our analysis to the genParticles, the MC-truth particle collection. The matching is performed for objects satisfying $\Delta R<0.4$ and a relative $p_{T}$ difference $|p_{T}^{(2)}-p_{T}^{(1)}|/|p_{T}^{(2)}|<3.0$, where object (2) is the RECO-object and (1) is the generator object. Additionally, we forbid two RECO objects from matching to the same generator-level object, and preference matches by minimising $\Delta R$. We then write the PDG IDs of the matched partons to our HDF5 files for use later.
Appendix B Comparison with prior work

### v2

@@ 4.3 Classification
In Fig. 9, we present both MC truth-derived (9a) and data-derived (9b, 9c, 9d) SIC curves for each classifier, shown as solid lines, and the substructure observables discussed in Secion 1, shown as dashed lines. Here, the label “TopicFlow” refers to a generative classifier of the form $p_{Q}(x)/p_{G}(x)$. The three data-derived panels assume different $f_{\mathrm{DoubleMu}},f_{\mathrm{Jet}}$ ($S$, $T$, $R$) as given in Tab. 4 . For the neural-network classifiers, curves are produced by averaging predictions of 10 independent models on the test set of the relevant dataset. Curves for the substructure observables are measured over the full dataset. The corresponding AUC scores are summarised in Tab. 5.
Figure 10: SIC curves constructed using samples from TopicFlow quark and gluon models. Each panel differs by the assumed quark fractions $f_{\mathrm{DoubleMu}},f_{\mathrm{Jet}}$ used to train the TopicFlow model, either LABEL:sub@fig:d3_topicflow_parton_roc $S$, LABEL:sub@fig:d3_topicflow_cwola_roc $T$ or LABEL:sub@fig:d3_topicflow_cwola_corrected_roc $R$ . Error bands show the standard deviation across 100 ($\times10$) TopicFlow/classifier pairs.
When evaluating classifiers on MC labels, as for Fig. 9a, the fully-supervised classifier is best, closely followed by the MC-CWoLa classifier. Save for jet mass, all of the substructure observables outperform the classifiers trained on data (Data CWoLa and TopicFlow). Considering next the data SIC curves of Figs 9b, 9c and 9d, the rankings are different. In particular, the Data CWoLa classifier achieves the maximum significance improvement while the fully-supervised network is bested by the $\sigma_{1}$ observable. [^8] Note that the peaks at low $\varepsilon_{Q}$ are likely artefacts due to low statistics in the testing set. TopicFlow and MC CWoLa achieve similar results to one another and have the next best performance after Data CWoLa. Comparing these data-derived panels to one another, we see that the assumed quark fractions $f_{\mathrm{DoubleMu}}$ and $f_{\mathrm{Jet}}$ do not impact the rankings among the classifiers. What does change between these two constructions is the absolute significance improvement of the models as well as the signal efficiency at which it is achieved. Specifically, the maximum SIC is highest in Fig. 9c which uses fractions that assume quarks and gluons are mutually irreducible in Data CWoLa predictions ($T$). On the other hand, the SIC curves in Fig. 9c, which use fractions extracted by Data CWoLa assuming the same $\kappa_{QG}$ as MC, have the lowest maxima. In fact, both panels in the right column indicate that little-to-no significance improvement can be yielded by any cut on the fully-supervised networks. Notably, no classifier in any of the data-derived panels reaches the maximum SIC achieved by the fully-supervised classifier in MC (Fig. 9a).
Finally, Fig. 10 shows SIC curves of each discriminative classifier as estimated using samples from TopicFlow quark and gluon distributions. [^9] TopicFlow generative classifiers are excluded since they represent the precise likelihood ratio for their samples and are therefore biased. We show three panels corresponding to choices $S$, $T$ or $R$ for the fractions $f_{\mathrm{DoubleMu}}$ and $f_{\mathrm{Jet}}$. The error bands show one standard deviation across the collection of curves produced by evaluating 10 classifiers on samples from 10 TopicFlow models. These bands therefore capture uncertainty in the TopicFlow model, which we find comprises most of the error. The general shapes of the curves agree with Fig. 9 except at low $\varepsilon_{Q}$ where the SIC for Data CWoLa previously exhibited large statistical variations. When amplifying the statistics with TopicFlow generated samples, the curves for each classifier are smooth and clearly exhibit a maximum. The TopicFlow SIC curves also maintain the same hierarchy among the three classifiers as Fig. 9. We note that the width of the bands on each curve are smaller than the differences between panels in Figs. 10b and 10c. Thus, the largest source of uncertainty in the expected discrimination power for a given quark gluon tagger in our datasets comes from the definition of the classes themselves. Again, as one expects, the greatest peak SIC is predicted by assuming mutually irreducible quarks and gluons since these are by definition maximally-separable categories.
5 Conclusions
Machine-learning methods that operate directly on real data offer a way to avoid test-time performance loss caused by imprecision of Monte Carlo simulation. In this work we have studied the application of weakly-supervised neural networks to quark/gluon classification using jets in the CMS Open Data. In particular, we took Z+Z+jet and dijet events as respective quark- and gluon- enriched samples to train CWoLa and TopicFlow models. The performance of these models was compared to fully-supervised networks trained on MC simulation.
In addition to the predictions based on parton labels in simulation, we made two estimates for the quark fractions of each real dataset. The first assumes that quarks and gluons are irreducible in the output of the classifiers, in line with the jet-topic framework. The second estimate assumes non-zero reducibility between quarks and gluons under the classifiers, as determined by MC. The results of both estimates differ from the MC prediction. These discrepancies may be covered by systematics that we cannot quantify in the CMS Open Data.
Using the extracted fractions, TopicFlow networks were trained to model the underlying quark and gluon distributions in real data. We employed a variant of the loss function introduced in Ref. Dolan:2022ikg that allows for more stable training due to a convex optimisation objective. The distributions encoded by these TopicFlow models fit the corresponding topics in data better than predictions from simulation—both pure quark/gluon categories and topics constructed in MC.
An interesting direction for future work would be to relax the assumption of sample independence of quarks and gluons in the mixed datasets. While we observed that sample independence is approximately satisfied in the Monte Carlo, we did not attempt to quantify the effects of its violation. As such, our results are subject to this extra degree of uncertainty. Some ideas to this end are given in Ref. Komiske:2018vkc . A different approach may be to match the two mixed distributions via reweighting as in Ref. ATLAS:2023pdx (and implicitly in Ref. Komiske:2022vxg ). Alternatively, techniques that directly morph between two distributions could be used, such as in Refs. Algren:2023qnb ; Golling:2023mqx ; Diefenbacher:2023flw ; Butter:2023ira . Beyond cases where sample dependence is only approximately satisfied, a dedicated treatment of this problem could extend the reach of weakly-supervised approaches to domains where the effect is known to be severe Lee:2023tfx .
Appendix A The CMS Open Data
A.1 Accessing the Open Data
The data from the 2010–2012 period can be accessed remotely or downloaded in the Analysis Object Data (AOD) format, a ROOT-compatible collection of data processed through the CMS software framework CMSSW CMSSW . The AODs are grouped into primary datasets on the basis of High-Level Trigger (HLT) paths that share common physics content. The typical size of an AOD file is a few gigabytes, and some thousands of AOD files constitute a primary dataset, which typically has a total size on the order of terabytes. Since we are interested in sourcing data samples with enriched quark- and gluon-jet content, our study makes use of the Jet jet2011A and DoubleMu doublemu2011A primary datasets, from which we extract gluon-rich dijet events and quark-rich $Z$ + jets events, respectively.
Our study of the 2011 Open Data takes place within version 5.3.32 of CMSSW, the recommended version compatible with the 2011 data. This software framework is used in data taking, processing, analysis and the generation of simulated samples. It defines the event model used to store the data and contains services and plugins we use in the initial stages of our data processing. The code is open source and available on GitHub CMSSW .
The CMSSW framework is also available through the CernVM file system (CVMFS) blomer_jakob_2020_4114078 . The open data release is accompanied by a virtual machine (VM) image, to which CVMFS is mounted, and this provides a simple interface to access the data and conduct analysis. Rather than using the VM, we instead use a light-weight container image executed on a batch computing system. The image contains the base operating system needed to execute CMSSW. This, along with a CVMFS mount, allows us to run the CMS software framework with only the data and some additional resources we discuss below.
Our approach to accessing the data is to download each AOD file on demand through the XRootD protocol xrootd , pass the events through our selection and analysis code within CMSSW, and write out much smaller, custom data files in the HDF5 format for downstream processing. This workflow allows us to work with multiple primary datasets without keeping additional terabytes of information we do not require in our study.
Our experience using the CMS Open Data has greatly benefited from the detailed former studies using the CMS Open Data and the increasing number of example analyses provided by CMS through the cms-opendata-validation repositories open-data-validation . Specifically, we have drawn on example code for the production of flat jet-tuples using the 2011 data 2011-jet-inclusivecrosssection-ntupleproduction and the general procedures of Ref. Tripathee:2017ybi and the MODProducer mod , adapted to work with the prescaled dijet triggers of the 2011 Jet Primary Dataset. We point the interested reader to Refs. Tripathee:2017ybi ; challenges_and_directions ; Bellis:2022onc for a detailed discussion of challenges and advice for using the CMS Open Data.
A.2 Analysing the Open Data
In addition to the AOD files, our analysis code requires access to the CMS condition database Guida_2015 ; conditiondatabase . This is a store of auxiliary, non-event-related data provided in the form of SQLite files within a local /cvmfs/cms-opendata-conddb.cern.ch directory inside the container image. In the context of our analysis these provide access to trigger information and to the jet energy corrections (JECs): MC-derived correction factors that mitigate detector noise and pile up. The condition database contains information relevant for different data-taking periods, each labelled by a global tag. Our analysis uses the global tags START53_LV6A1::All for simulation and FT_53_LV5_AN1::All for data. A connection is established to the conditions database within our configuration file.
The jet constituents are the objects obtained from the Particle Flow (PF) algorithm CMS-PAS-PFT-09-001 ; CMS-PAS-PFT-10-001 ; CMS:2017yfk . In our analysis code, we pull jets from the ak5PFJets collection, corresponding to Fastjet Cacciari:2011ma anti-$k_{t}$ jets Cacciari:2008gp reconstructed from the PF particles with $R=0.5$. The constituent and jet $p_{T}$, along with the jet energy fractions (used later to implement our jet-quality requirements), are scaled by the JECs mentioned earlier. These are sourced using the ak5PFL1FastL2L3Residual correction label for data, and ak5PFL1FastL2L3 for MC.
Another layer of processing is required to deal with pileup: contamination from additional, non-event vertices. Here, we apply the procedure of Charged Hadron Subtraction (CHS) CMS:2014ata , where tracks not associated with the event vertex (defined below) are pruned from a jet. We draw vertex information from the standard vertex collection provided by CMSSW within our configuration file with the requirements that the minimum number of degrees of freedom in the vertex fit be larger than 4; $|z|\leq 24$ mm, where $z$ is the $z$-coordinate of the point of closest approach of the tracks to the $z$-axis; and $\rho<2$, where $\rho$ is an estimate of the average amount of transverse energy in the event per unit area coming from soft activity such as pileup. From among these vertices, the event vertex is defined as that primary vertex whose tracks have the largest sum of squared $p_{T}$, while all other interaction vertices are classed as pileup vertices. Within the EDAnalyzer we write out the vertex-fit information associated with the charged hadrons and later remove those assigned to pileup vertices. The distribution of the number of offline primary vertices (NPV) looks different in simulation compared to data for the triggers of interest to our study. To correct this, we apply an additional weight to the MC that makes the NPV distributions match.
The remaining PF candidates are used to calculate the jet observables that form the basis of our machine-learning study. (These are introduced in more detail in Sec. 3.) These observables are IRC-safe, and therefore an additional cut on the $p_{T}$ of the jet-constituent objects is unnecessary.
The muon objects that enter our analysis are required to pass the GlobalMuonPromptTight requirements. These include that the track should be identified as a global muon CMS:2012nsv ; CMS:2018rym , the associated normalised $\chi^{2}$ in the global muon fit should be less than 10, and the number of valid muon-detector hits used in the fit should be non-zero.
In order to associate truth-level parton information with the PF candidates in MC, the generator objects must be matched to the PF objects from offline reconstruction (RECO) through MC matching. We follow the implementation of this matching from the CMSSW Physics Analysis Toolkit Adam:1196152 . Specifically, we call the MCMatcher EDProducer within our configuration file. This matches jets from the ak5PFJets collection we use in our analysis to the genParticles, the MC-truth particle collection. The matching is performed for objects satisfying $\Delta R<0.4$ and a relative $p_{T}$ difference $|p_{T}^{(2)}-p_{T}^{(1)}|/|p_{T}^{(2)}|<3.0$, where object (2) is the RECO-object and (1) is the generator object. Additionally, we forbid two RECO objects from matching to the same generator-level object, and preference matches by minimising $\Delta R$. We then write the PDG IDs of the matched partons to our HDF5 files for use later.
Appendix B Definitions of substructure observables
Energy flow polynomials (EFPs)
where $M$ is the number of jet constituents, $z_{a}=p_{T,a}/p_{T,J}$ is the $p_{T}$ fraction of constituent $a$ relative to the jet, $\theta_{ab}$ is the separation of $a$ and $b$ in the rapidity-azimuth plane, and $\beta$ is a parameter. EFPs are manifestly IRC safe.
Jet fragmentation distribution $p_{T}^{D}$
The jet fragmentation distribution is defined as
@@ Appendix B Definitions of substructure observables
Jet major axis $\sigma_{1}$
The major axis of a jet is determined by calculating the eigenvalues of the matrix
@@
Appendix C Effects from parton matching
In this appendix, we study the effect of removing non-parton-matched jets from the simulated samples.
Figure 11: Histograms of LABEL:sub@fig:allmatch_preds the MC CWoLa network output LABEL:sub@fig:allmatch_sigma_1 $\sigma_{1}$ and LABEL:sub@fig:allmatch_ptD $p_{T}^{D}$ for matched and unmatched jets in QCDSim and DYSim MC.
Figure 11 shows the distributions of jet substructure observables for the leading jet in QCD and DY MC, split by whether or not the jet has a parton match. We show standard quark/gluon discriminators $\sigma_{1}$ and $p_{T}^{D}$ as well as the MC CWoLa classifier score.
Based on these variables, we identify minimal difference between matched and unmatched jets in QCDSim. On the other hand, we see that the unmatched jets in DYSim behave as jets in QCDSim. This suggests that the unmatched DYSim sample is enriched in gluon-initiated jets. Consequently, the gluon fraction of the full DYSim sample is likely higher than the estimate based on particle IDs. However, by assuming that the quark fraction of unmatched jets in DYSim is 30% (i.e. equal to the QCDSim sample) and taking into account that only 3% of DYSim jets are unmatched, we expect the true quark fraction in DYSim to be 0.727, just 2% smaller than our nominal value from Eq. 12.
We can also check how the classifier performance is affected by training MC CWoLa networks (i.e. DYSim vs QCDSim) including or excluding jets without parton matches. Table 6 shows the AUC score for networks trained and evaluated on different subsets of the simulated datasets. For training, we consider either all jets, or only those with a parton match. For testing, we calculate the AUC separately over jets with a match and jets without. In all cases, the AUCs agree within errors, which we estimate as the standard deviation over 30 networks. We therefore conclude that the removal of parton-matched jets from the simulated samples does not have a large effect on the classifier results.
Train set
Matched All
Test set Matched $7\pm 0.002 0.606±0.0$$6\pm 0.002$
Unmatched $0.607\pm 0.002$ $0.605\pm 0.002$
Table 6: AUC scores for MC CWoLa networks (DYSim vs QCDSim) trained with or without non-parton-matched jets, and evaluated over matched and non-matched jets separately. Errors are the standard deviation over 30 networks.
Appendix D Comparison with prior work